# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v5

@@ 2.2 Text-to-Image Model
Ming-Omni [171] b-3 Ling QwenViT Learnable Query Multi-scale DiT Causal 2025-06
@@ 3.1 Diffusion Models
Building upon this foundation, MMaDA [129] scales up the diffusion paradigm toward a unified multimodal foundation model. It adopts LLaDA-8B-Instruct [212] as the language backbone and uses a MAGVIT-v2 [213] image tokenizer to convert images into discrete semantic tokens. This unified token space enables seamless multimodal conditioning during generation. To improve alignment across modalities, MMaDA introduces a mixed chain-of-thought (CoT) fine-tuning strategy, which unifies reasoning formats between text and vision tasks. This alignment facilitates cold-start reinforcement learning, allowing effective post-training from the outset. Furthermore, MMaDA incorporates a novel UniGRPO method, a unified policy-gradient-based RL algorithm designed for diffusion models. UniGRPO enables post-training optimization across both reasoning and generation tasks by leveraging diversified reward signals, such as factual correctness, visual-textual alignment, and user preferences. This design ensures the model consistently improves across a broad range of capabilities, rather than overfitting to a narrow task-specific reward.
@@ 3.2 Auto-Regressive Models
Recent works have adopted and enhanced pixel-based tokenization with various encoder designs. LWM [29] employs a VQGAN tokenizer [32] to encode images into discrete latent codes without requiring semantic supervision. It proposes a multimodal world modeling framework, wherein visual and textual tokens are serialized together for unified autoregressive modeling. By learning world dynamics purely through reconstruction-based visual tokens and textual descriptions, LWM demonstrates that large-scale multimodal generation is feasible without specialized semantic tokenization. Both Chameleon [30] and ANOLE [132] adopt VQ-IMG [222], an improved VQ-VAE variant designed for content-rich image generation. Compared to standard VQGAN tokenizers, VQ-IMG features a deeper encoder with larger receptive fields and incorporates residual prediction to better preserve complex visual details. This enhancement enables Chameleon and ANOLE to serialize image content more faithfully, thereby supporting high-quality multimodal generation. Moreover, these models facilitate interleaved generation, allowing text and image tokens to be generated alternately within a unified autoregressive framework. Emu3 [133], SynerGen-VL [136] and UGen [138] employs SBER-MoVQGAN [221, 220], a multi-scale VQGAN variant that encodes images into latent representations capturing both global structure and fine-grained details. By leveraging multi-scale tokenization, these models improve the expressiveness of visual representations for autoregressive modeling while maintaining efficient training throughput. Similar with LWM [29], Liquid [137] utilizes a VQGAN-style tokenizer and uncovers a novel insight that visual understanding and generation can mutually benefit when unified under a single autoregressive objective and shared visual token representation. Moreover, MMAR [134], Orthus [135], Harmon [139] introduce the frameworks that utilize continuous-valued image tokens extracted by their corresponding encoders, avoiding the information loss associated with discretization. They also decouple the diffusion process from the AR backbone by employing lightweight diffusion heads atop each auto-regressed image patch embedding. This design ensures that the backbone’s hidden representations are not confined to the final denoising step, facilitating better image understanding. TokLIP [140] integrates a low-level discrete VQGAN tokenizer with a ViT-based token encoder SigLIP [209] to capture high-level continuous semantics, which not only empowers visual tokens with high-level semantic understanding but also enhances low-level generative capacity. Selftok [141] introduces a novel discrete visual self-consistency tokenizer, achieving a favorable trade-off between high-quality reconstruction and compression rate while enabling optimal policy improvement for effective visual reinforcement learning.
@@
Across these models except for MMAR [134] and Harmon [139], causal attention masks are applied during both pretraining and generation phases, ensuring that each token only attends to preceding tokens in the sequence. They are trained using a next-token prediction loss, where both image and text tokens are predicted autoregressively, thus unifying the training objective across modalities. Notably, in pixel-based encoding approaches, the decoder used to reconstruct images from latent tokens typically follows the paired decoder structure originally proposed in VQGAN-like models. These decoders are lightweight convolutional architectures specifically optimized to map discrete latent grids back to the pixel space, focusing primarily on accurate low-level reconstruction rather than high-level semantic reasoning. Moreover, since some methods, like MMAR [134], Orthus [135] and Harmon [139], tokenize the image into continuous latents, they adopt the lightweight diffusion MLP as their decoder to map continuous latents back to the pixel space.
@@
2) Semantic Encoding. To overcome the semantic limitations inherent in pixel-based encoders, a growing body of work adopts semantic encoding, where image inputs are processed using pretrained text-aligned vision encoders such as OpenAI-CLIP [22], SigLIP [209], EVA-CLIP [36], or more recent unified tokenizers like UNIT [223], as shown in Fig. 5 (b-2). Some of these models leverage the multimodal features encoded by the multimodal autoregressive model as conditions for a diffusion model, enabling image generation with retained multimodal understanding capabilities, like OmniGen2 [158] that utilizes Qwen2.5-VL [10] as the multimodal model and enhanced OmniGen [224] as the image diffusion model, Ovis-U1 [159] extending the multimodal model Ovis [12] into a unified model by incorporating a custom-designed diffusion transformer while Qwen-Image [161] similarly building upon Qwen2.5-VL [10] by integrating a diffusion transformer. However, most of these models are trained on large-scale image-text pairs with contrastive or regression-based objectives, producing visual embeddings that align closely with language features in a shared semantic space. Such representations enable more effective cross-modal alignment and are particularly beneficial for multimodal understanding and generation.
@@
Several representative models leverage different semantic encoders and architectural designs to support unified multimodal tasks. Emu [142], Emu2 [33], and LaViT [143] all employ EVA-CLIP [36] as their vision encoder. Notably, Emu [142] introduces the initial architecture combining a frozen EVA-CLIP encoder, a large language model, and a diffusion decoder to unify VQA, image captioning, and image generation. Emu2 [33] builds upon Emu [142] by proposing a simplified and scalable modeling framework for unified multimodal pretraining. It scales the MLLM model up to 37B parameters, significantly enhancing both understanding and generation capabilities. Bifrost-1 [162] employs two semantic encoders, ViT for generation and the one employed in the used MLLM (QWen2.5-VL) for understanding. The predicted CLIP latents are used to bridge the MLLM and diffusion model. LaViT [143] introduces a dynamic visual tokenization mechanism built on top of EVA-CLIP. It employs a selector and merger module to adaptively select visual tokens from image embeddings based on content complexity. This process dynamically determines the length of the visual token sequence per image. The dynamic tokenization significantly reduces redundant information while preserving important visual cues, improving training efficiency and generation quality in tasks such as captioning, visual question answering, and image generation. DreamLLM [34], VL-GPT [35], and MM-Interleaved [144], and PUMA [147] utilize OpenAI-CLIP encoder [22]. DreamLLM [34] introduces a lightweight linear projection to align CLIP embeddings with language tokens, while VL-GPT [35] employs a powerful casual transformer after OpenAI-CLIP vision encoder to effectively retain both semantic information and pixel details of the original image. Both MM-Interleaved [144] and PUMA [147] extract multi-granular image features via a CLIP tokenizer with simple ViT-Adapter or pooling operation to provide fine-grained feature fusion, thus supporting rich multimodal generation. Mini-Gemini [145] introduces a visual token enhancement mechanism that requires dual semantic encoders. Specifically, it leverages a CLIP-pretrained ViT encoder [22] to obtain global visual tokens, while a LAION-pretrained ConvNeXt encoder provides dense local visual information. A cross-attention module is then employed to refine the global visual tokens by incorporating detailed visual cues from the dense encoder. These enhanced global tokens are subsequently combined with text tokens and processed by an LLM for joint vision-language understanding and generation. This design effectively bridges the semantic abstraction of CLIP features with the pixel-level precision of dense encoders. MetaMorph [148] employs SigLIP [209] to extract visual embeddings and introduces modality-specific adapters within a pretrained language model. These adapters are inserted throughout multiple transformer layers, allowing for deeper vision-language interaction compared to shallow projection approaches. ILLUME [149] adopt UNIT [223] as its vision encoder to provide a unified representation that balances semantic alignment and pixel-level fidelity. Unlike CLIP-like encoders that focus purely on contrastive objectives, UNIT [223] is jointly trained with both image reconstruction and contrastive alignment losses, producing tokens suitable for both vision-language understanding and image synthesis. Built on the powerful UNIT tokenizer, ILLUME effectively generates image tokens that retain both semantic and pixel-level information, which achieves better performance in multiple understanding and generation tasks, including captioning, VQA, text-to-image, and interleaved generation. Similarly, VILA-U [146] and UniTok [150] mimic UNIT [223] to introduce image-text contrastive learning to obtain a novel text-aligned vision tokenizer that balances semantic alignment and pixel-level fidelity. QLIP [151] addresses the potential conflict between reconstruction and text-image alignment tasks by implementing binary-spherical quantization. Tar [157] initiates the visual codebook by leveraging the vocabulary of LLMs and incorporates scale-adaptive pooling and decoding methodologies. This approach enables the model to adjust the length of the tokenizer according to the requirement: employing coarse-grained tokenizers for efficient generation and fine-grained tokenizers for comprehensive understanding. In generation tasks, Tar utilizes diffusion techniques to enhance the visual generation outcomes of AR models. UniFork [153] capitalizes on the text-aligned vision features of VILA-U. However, differentiating itself from the fully-shared parameters of understanding and generation MLLM, UniFork shares the parameters solely with these tasks at the shallow layer. At the deeper layer, these tasks are managed by distinct networks. This architecture successfully mediates the equilibrium between shared learning and task-specific specialization. UniCode2 [154] employs a cascaded codebook. In line with the method outlined in [225], it utilizes a substantial codebook derived from clustered SigLIP feature as the frozen foundational codebook, while introducing supplementary learnable codebooks to refine semantics specific to particular tasks. This separation enhances utilization and fosters robust learning. Recent work DualToken [152] uses shallow-layer features of SigLIP for reconstruction and deep-layer features of SigLIP for semantic learning, thereby obtaining the texture and semantic visual features simultaneously. As a result, DualToken [152] achieves superior performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks. X-Omni [160] utilizes SigLIP-VQ as a visual encoder and employs reinforcement learning to mitigate the cumulative error associated with autoregressive inference and to reduce the information loss inherent in discrete encoding. This methodology substantially enhances the generation quality of discrete autoregressive models, facilitating a seamless integration of image and language generation.
@@
Across most of these models, causal attention masks are applied during MLLM training, and next-token prediction loss is used to optimize both text and vision token generation. For image generation, most of these models typically employ diffusion-based decoders, such as SD family [14, 226], IP-adapter [227], FLUX [16], and Lumina-Next [228], which are trained independently from the MLLM. During inference, the MLLM produces semantic-level visual tokens, which are then passed to the diffusion decoder for final image synthesis. This design choice—pairing semantic encoders with diffusion decoders—is motivated by the fact that semantic embeddings encode high-level conceptual information but lack the spatial density and low-level granularity required for direct pixel reconstruction. Diffusion models, with their iterative denoising mechanisms, are particularly well-suited for this setting: they are capable of progressively refining semantic representations into high-resolution, photorealistic images, even when the input tokens are sparse or abstract. In contrast, although few approaches (i.e., VILA-U [146] and UniTok [150]) adopt pixel-based decoders, their generated image quality is less competitive than the diffusion decoders. Thus, diffusion decoders provide a more robust and expressive decoding pathway for semantically compressed visual tokens, significantly improving text-image alignment, global coherence, and visual fidelity. UniWorld [155] and Pisces [156] have endeavored to develop and expand such a solution. UniWorld directly utilizes the output features of pre-trained MLLM for visual comprehension as a high-level conditional signal, while employing SigLIP as low-level conditional signal to deliver comprehensive semantic visual control for DiT. Pisces employs EVA-CLIP as a condition for visual generation tasks and leverages diffusion to further enhance the model’s visual generation output. For various tasks, Pisces introduces tailored visual vector lengths and employs distinct MLPs to encode conditions. This approach increases the flexibility of model design while mitigating the inference cost compared to a single encoder configuration.
@@
Current implementations of learnable query encoding can be broadly divided into two representative paradigms. The first is represented by SEED [163], which proposes a seed tokenizer that learns causal visual embeddings. Specifically, an input image is first encoded into dense token features via a BLIP-2 ViT encoder [53]. These features are then concatenated with a set of learnable query tokens and processed by a causal Q-Former to produce causal visual embeddings. This design is trained using both image-text contrastive learning and image reconstruction supervision, allowing the learned embeddings to simultaneously retain low-level visual detail and capture high-level semantic alignment with text. Building on this foundation, SEED-LLAMA [164] and SEED-X [165] enhance the model’s capacity by replacing the OPT backbone [229] with a stronger LLaMA2 model [2] and upgrading the decoder to UnCLIP-SD [14] or SDXL [226], leading to improved performance in both understanding and generation tasks. The second approach, introduced by MetaQueries [166], provides a simplified version of learnable query encoding. Here, image features are extracted via a frozen SigLIP encoder [209], which are then concatenated with learnable query tokens and directly passed through a frozen vision-language backbone such as LLaVA [216] or Qwen2.5-VL [10]. The output causal embeddings are used as conditioning inputs for a diffusion-based image decoder, enabling high-quality image generation. Because the backbone is kept frozen, the vision-language understanding capabilities remain consistent with the underlying pretrained models, offering a lightweight yet effective solution for multimodal generation. Open-Uni [170] refines the architecture of MetaQueries by utilizing solely learnable queries and a lightweight connector between a MLLM and a diffusion model, facilitating cohesive multimodal understanding and generation. Open-Uni demonstrates that the connetor between the MLLM visual understanding component and the diffusion-based visual generation component can be minimal in complexity, exemplified by a configuration comprising merely six Transformer layers. Nexus-Gen [167] and Ming-Lite-Uni [168] follow the MetaQueries paradigm, but with notable advancements to further enhance multimodal generation. Nexus-Gen [167] introduces a more powerful diffusion decoder, FLUX-1.dev, which significantly improves the generation quality. This approach allows the model to better capture the intricate details and high-fidelity features necessary for complex image generation tasks. On the other hand, Ming-Lite-Uni [168] takes a different route by introducing a highly capable MLLM model, M2-omini [200], for enhanced vision-language interaction. This model performs advanced vision-language conditioning to generate the conditioned image embeddings, ensuring a more semantically aligned representation. In addition, Ming-Lite-Uni fine-tunes its diffusion model by incorporating multi-scale learnable tokens, which facilitate improved semantic alignment across various visual scales. The multi-scale representation alignment mechanism enhances the model’s ability to generate detailed and contextually rich images from textual prompts, addressing challenges such as resolution mismatches and semantic inconsistencies. This innovative approach makes Ming-Lite-Uni a powerful tool for multimodal understanding and generation, pushing the boundaries of current methods in both flexibility and performance. Ming-Omni [171] adheres to the integrated MoE architecture, wherein modality-specific routing is facilitated through dedicated mechanisms tailored for each token, thereby enabling customized routing distributions. To address the multi-scale phenomenon inherent in visual generation [113], Ming-Omni employs multi-scale learnable queries, directed by an alignment strategy, to iteratively generate images progressing from coarse to fine detail. Furthermore, Ming-Omni integrates the audio modality and implements a dual-stage training strategy to mitigate the mutual influence between audio comprehension and generation tasks. The initial stage emphasizes comprehension capabilities, while the subsequent stage concentrates on enhancing generation quality. BLIP3-o [169] also employs learnable queries to bridge multimodal understanding and generation. However, it utilizes two diffusion models: one for learning CLIP embeddings and the other for using CLIP as a condition to generate images. It reveals that flow matching loss is more effective than MSE loss, enabling more diverse image sampling and yielding better image quality. UniLIP [172] incrementally incorporates reconstruction ability into CLIP through the self-distillation, then employs a learnable query along with the hidden state of the last layer of the MLLM as combined conditions. This framework is demonstrated to optimize the abundant information for visual editing. To exploit the hierarchical representations within the MLLM’s intermediate layers, TBAC-UniImage [173] applies learnable queries at multiple layers instead of the last layer. To sum up, these learnable query-based designs share a common strength: they provide adaptive, compact, and semantically enriched representations that support both efficient image understanding and high-quality generation. By focusing on task-driven token extraction, such models offer a flexible and extensible alternative to traditional visual tokenizers, especially in unified multimodal frameworks.

### v6

@@ 2.2 Text-to-Image Model
Lavida-O [161] a LaViDa SigLIP VQ-Encoder VQ-Encoder Bidirect. 2025-09
UniModel [373] a MMDiT in Qwen-Image Wan-2.1-VAE Wan-2.1-VAE Bidirect. 2025-11
@@
OneCat [152] b-1 Qwen2.5 Convolution + MLP VAE in Infinity VAE in Infinity Causal 2025-09
Emu3.5 [307] b-1 Modified Qwen3 SBER-MoVQGAN SBER-MoVQGAN Causal 2025-10
@@
Ming-UniVision [123] b-2 Ming-UniVision MingTok MingTok Causal 2025-10
MammothModa2 [252] b-2 Qwen3-VL-8B QwenViT MammothTok Single-stream DiT Causal 2025-11
@@
UniPic 2.0 [317] b-3 Qwen2.5-VL QwenViT Learnable Query SD3.5-Medium Causal 2025-09
@@
LightFusion [311] c-2 QWen2.5-VL +Wan2.2-TI2V QWen2.5-VL Wan2.2-TI2V DCAE Bidirect. 2025-11
HBridge [305] c-2 QWen + OmniGen2 QwenViT SigLip OmniGen Bidirect. 2025-11
EMMA [105] c-2 Qwen3-34B SigLIP DCAE DCAE Bidirect. 2025-12
@@ 3.1 Diffusion Models
Building upon this foundation, MMaDA [343] scales up the diffusion paradigm toward a unified multimodal foundation model. It adopts LLaDA-8B-Instruct [209] as the language backbone and uses a MAGVIT-v2 [358] image tokenizer to convert images into discrete semantic tokens. This unified token space enables seamless multimodal conditioning during generation. To improve alignment across modalities, MMaDA introduces a mixed chain-of-thought (CoT) fine-tuning strategy, which unifies reasoning formats between text and vision tasks. This alignment facilitates cold-start reinforcement learning, allowing effective post-training from the outset. Furthermore, MMaDA incorporates a novel UniGRPO method, a unified policy-gradient-based RL algorithm designed for diffusion models. UniGRPO enables post-training optimization across both reasoning and generation tasks by leveraging diversified reward signals, such as factual correctness, visual-textual alignment, and user preferences. This design ensures the model consistently improves across a broad range of capabilities, rather than overfitting to a narrow task-specific reward. Built on LaViDa [162], Lavida-O [161] bridges the gap between unified multi-modal diffusion models and state-of-the-art AR models by implementing a resource-efficient Elastic Mixture-of-Transformers (Elastic-MoT) architecture. To overcome the scarcity of effective masked generative training, it employs progressive upscaling and token compression for robust model scaling. Furthermore, it distinguishes itself through planning and self-reflection mechanisms, which allow the model to use its understanding capability to actively refine and improve generation outputs.
Building upon the insight from DeepSeek-OCR [316] that text can be processed as a visual signal, UniModel [373] unifies these tasks by rendering text into images and employing a diffusion model for both understanding and generation. While this represents a novel strategy, significant challenges remain, including text rendering fidelity, multilingual generation capabilities, and long-context modeling.
@@ 3.2 Auto-Regressive Models
Recent works have adopted and enhanced pixel-based tokenization with various encoder designs. LWM [183] employs a VQGAN tokenizer [72] to encode images into discrete latent codes without requiring semantic supervision. It proposes a multimodal world modeling framework, wherein visual and textual tokens are serialized together for unified autoregressive modeling. By learning world dynamics purely through reconstruction-based visual tokens and textual descriptions, LWM demonstrates that large-scale multimodal generation is feasible without specialized semantic tokenization. Both Chameleon [276] and ANOLE [51] adopt VQ-IMG [81], an improved VQ-VAE variant designed for content-rich image generation. Compared to standard VQGAN tokenizers, VQ-IMG features a deeper encoder with larger receptive fields and incorporates residual prediction to better preserve complex visual details. This enhancement enables Chameleon and ANOLE to serialize image content more faithfully, thereby supporting high-quality multimodal generation. Moreover, these models facilitate interleaved generation, allowing text and image tokens to be generated alternately within a unified autoregressive framework. Emu3 [307], SynerGen-VL [153] and UGen [275] employs SBER-MoVQGAN [236, 389], a multi-scale VQGAN variant that encodes images into latent representations capturing both global structure and fine-grained details. By leveraging multi-scale tokenization, these models improve the expressiveness of visual representations for autoregressive modeling while maintaining efficient training throughput. EMU3.5 [58] develops Discrete Diffusion Adaptation to accelerate inference via bidirectional parallel prediction. Similar with LWM [183], Liquid [321] utilizes a VQGAN-style tokenizer and uncovers a novel insight that visual understanding and generation can mutually benefit when unified under a single autoregressive objective and shared visual token representation. Moreover, MMAR [342], Orthus [140], Harmon [324] introduce the frameworks that utilize continuous-valued image tokens extracted by their corresponding encoders, avoiding the information loss associated with discretization. They also decouple the diffusion process from the AR backbone by employing lightweight diffusion heads atop each auto-regressed image patch embedding. This design ensures that the backbone’s hidden representations are not confined to the final denoising step, facilitating better image understanding. TokLIP [180] integrates a low-level discrete VQGAN tokenizer with a ViT-based token encoder SigLIP [371] to capture high-level continuous semantics, which not only empowers visual tokens with high-level semantic understanding but also enhances low-level generative capacity. Selftok [295] introduces a novel discrete visual self-consistency tokenizer, achieving a favorable trade-off between high-quality reconstruction and compression rate while enabling optimal policy improvement for effective visual reinforcement learning. OneCat [152] employs a patch embedding layer to map raw images into continuous tokens for understanding and editing, alongside a pre-trained multi-scale VAE [102] for reconstruction. Notably, it adopts a hybrid generation strategy: standard next-token prediction for text and next-scale prediction [102] for image synthesis.
@@
Across most of these models, causal attention masks are applied during both pretraining and generation phases, ensuring that each token only attends to preceding tokens in the sequence. They are trained using a next-token prediction loss, where both image and text tokens are predicted autoregressively, thus unifying the training objective across modalities. Notably, in pixel-based encoding approaches, the decoder used to reconstruct images from latent tokens typically follows the paired decoder structure originally proposed in VQGAN-like models. These decoders are lightweight convolutional architectures specifically optimized to map discrete latent grids back to the pixel space, focusing primarily on accurate low-level reconstruction rather than high-level semantic reasoning. Moreover, since some methods, like MMAR [342], Orthus [140] and Harmon [324], tokenize the image into continuous latents, they adopt the lightweight diffusion MLP as their decoder to map continuous latents back to the pixel space.
@@
2) Semantic Encoding. To overcome the semantic limitations inherent in pixel-based encoders, a growing body of work adopts semantic encoding, where image inputs are processed using pretrained text-aligned vision encoders such as OpenAI-CLIP [230], SigLIP [371], EVA-CLIP [268], or more recent unified tokenizers like UNIT [399], as shown in Fig. 5 (b-2). Some of these models leverage the multimodal features encoded by the multimodal autoregressive model as conditions for a diffusion model, enabling image generation with retained multimodal understanding capabilities, like OmniGen2 [320] that utilizes Qwen2.5-VL [13] as the multimodal model and enhanced OmniGen [332] as the image diffusion model, Ovis-U1 [298] extending the multimodal model Ovis [199] into a unified model by incorporating a custom-designed diffusion transformer while Qwen-Image [318] similarly building upon Qwen2.5-VL [13] by integrating a diffusion transformer. UniVideo [314] and Omni-Video [273] extend this paradigm to the video modality. However, most of these models are trained on large-scale image-text pairs with contrastive or regression-based objectives, producing visual embeddings that align closely with language features in a shared semantic space. Such representations enable more effective cross-modal alignment and are particularly beneficial for multimodal understanding and generation.
@@
Several representative models leverage different semantic encoders and architectural designs to support unified multimodal tasks. Emu [269], Emu2 [267], and LaViT [135] all employ EVA-CLIP [268] as their vision encoder. Notably, Emu [269] introduces the initial architecture combining a frozen EVA-CLIP encoder, a large language model, and a diffusion decoder to unify VQA, image captioning, and image generation. Emu2 [267] builds upon Emu [269] by proposing a simplified and scalable modeling framework for unified multimodal pretraining. It scales the MLLM model up to 37B parameters, significantly enhancing both understanding and generation capabilities. Bifrost-1 [179] employs two semantic encoders, ViT for generation and the one employed in the used MLLM (QWen2.5-VL) for understanding. The predicted CLIP latents are used to bridge the MLLM and diffusion model. LaViT [135] introduces a dynamic visual tokenization mechanism built on top of EVA-CLIP. It employs a selector and merger module to adaptively select visual tokens from image embeddings based on content complexity. This process dynamically determines the length of the visual token sequence per image. The dynamic tokenization significantly reduces redundant information while preserving important visual cues, improving training efficiency and generation quality in tasks such as captioning, visual question answering, and image generation. DreamLLM [66], VL-GPT [395], and MM-Interleaved [283], and PUMA [75] utilize OpenAI-CLIP encoder [230]. DreamLLM [66] introduces a lightweight linear projection to align CLIP embeddings with language tokens, while VL-GPT [395] employs a powerful casual transformer after OpenAI-CLIP vision encoder to effectively retain both semantic information and pixel details of the original image. Both MM-Interleaved [283] and PUMA [75] extract multi-granular image features via a CLIP tokenizer with simple ViT-Adapter or pooling operation to provide fine-grained feature fusion, thus supporting rich multimodal generation. Mini-Gemini [169] introduces a visual token enhancement mechanism that requires dual semantic encoders. Specifically, it leverages a CLIP-pretrained ViT encoder [230] to obtain global visual tokens, while a LAION-pretrained ConvNeXt encoder provides dense local visual information. A cross-attention module is then employed to refine the global visual tokens by incorporating detailed visual cues from the dense encoder. These enhanced global tokens are subsequently combined with text tokens and processed by an LLM for joint vision-language understanding and generation. This design effectively bridges the semantic abstraction of CLIP features with the pixel-level precision of dense encoders. MetaMorph [286] employs SigLIP [371] to extract visual embeddings and introduces modality-specific adapters within a pretrained language model. These adapters are inserted throughout multiple transformer layers, allowing for deeper vision-language interaction compared to shallow projection approaches. ILLUME [297] adopt UNIT [399] as its vision encoder to provide a unified representation that balances semantic alignment and pixel-level fidelity. Unlike CLIP-like encoders that focus purely on contrastive objectives, UNIT [399] is jointly trained with both image reconstruction and contrastive alignment losses, producing tokens suitable for both vision-language understanding and image synthesis. Built on the powerful UNIT tokenizer, ILLUME effectively generates image tokens that retain both semantic and pixel-level information, which achieves better performance in multiple understanding and generation tasks, including captioning, VQA, text-to-image, and interleaved generation. Similarly, VILA-U [326] and UniTok [200] mimic UNIT [399] to introduce image-text contrastive learning to obtain a novel text-aligned vision tokenizer that balances semantic alignment and pixel-level fidelity. QLIP [387] addresses the potential conflict between reconstruction and text-image alignment tasks by implementing binary-spherical quantization. Tar [101] initiates the visual codebook by leveraging the vocabulary of LLMs and incorporates scale-adaptive pooling and decoding methodologies. This approach enables the model to adjust the length of the tokenizer according to the requirement: employing coarse-grained tokenizers for efficient generation and fine-grained tokenizers for comprehensive understanding. In generation tasks, Tar utilizes diffusion techniques to enhance the visual generation outcomes of AR models. UniFork [164] capitalizes on the text-aligned vision features of VILA-U. However, differentiating itself from the fully-shared parameters of understanding and generation MLLM, UniFork shares the parameters solely with these tasks at the shallow layer. At the deeper layer, these tasks are managed by distinct networks. This architecture successfully mediates the equilibrium between shared learning and task-specific specialization. UniCode2 [27] employs a cascaded codebook. In line with the method outlined in [396], it utilizes a substantial codebook derived from clustered SigLIP feature as the frozen foundational codebook, while introducing supplementary learnable codebooks to refine semantics specific to particular tasks. This separation enhances utilization and fosters robust learning. Recent work DualToken [262] uses shallow-layer features of SigLIP for reconstruction and deep-layer features of SigLIP for semantic learning, thereby obtaining the texture and semantic visual features simultaneously. As a result, DualToken [262] achieves superior performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks. X-Omni [87] utilizes SigLIP-VQ as a visual encoder and employs reinforcement learning to mitigate the cumulative error associated with autoregressive inference and to reduce the information loss inherent in discrete encoding. This methodology substantially enhances the generation quality of discrete autoregressive models, facilitating a seamless integration of image and language generation. Ming-UniVision [123] introduces MingTok, a tokenizer comprising three modules: a low-level encoder, a semantic decoder, and a pixel decoder. The low-level encoder projects the input image into a compact latent space, which is subsequently processed by the semantic decoder for understanding or editing tasks. For generation, predicted tokens within this latent space are reconstructed into raw images via the pixel decoder. MammothModa2 [252] introduces MammothTok, a unified visual tokenizer built upon AIMv2 for generation, while utilizing QwenViT for understanding. Notably, it leverages multi-layer features from the understanding component to condition the generation process. Recently, RAE [388] demonstrated that semantic encoders can also be well adapted to diffusion models, and thus using semantic encoders as visual tokenizers for unified models might become the future direction. VQRAE [68] has demonstrated that discretized SigLIP features are effective for both understanding and generation tasks. However, the potential for a fully unified framework based on these features remains underexplored..
@@
Across most of these models, causal attention masks are applied during MLLM training, and next-token prediction loss is used to optimize both text and vision token generation. For image generation, most of these models typically employ diffusion-based decoders, such as SD family [244, 224], IP-adapter [347], FLUX [143], and Lumina-Next [402], which are trained independently from the MLLM. During inference, the MLLM produces semantic-level visual tokens, which are then passed to the diffusion decoder for final image synthesis. This design choice—pairing semantic encoders with diffusion decoders—is motivated by the fact that semantic embeddings encode high-level conceptual information but lack the spatial density and low-level granularity required for direct pixel reconstruction. Diffusion models, with their iterative denoising mechanisms, are particularly well-suited for this setting: they are capable of progressively refining semantic representations into high-resolution, photorealistic images, even when the input tokens are sparse or abstract. In contrast, although few approaches (i.e., VILA-U [326], UniTok [200], and VQRAE [68]) adopt pixel-based decoders, their generated image quality is less competitive than the diffusion decoders. Thus, diffusion decoders provide a more robust and expressive decoding pathway for semantically compressed visual tokens, significantly improving text-image alignment, global coherence, and visual fidelity. UniWorld [178] and Pisces [340] have endeavored to develop and expand such a solution. UniWorld directly utilizes the output features of pre-trained MLLM for visual comprehension as a high-level conditional signal, while employing SigLIP as low-level conditional signal to deliver comprehensive semantic visual control for DiT. Pisces employs EVA-CLIP as a condition for visual generation tasks and leverages diffusion to further enhance the model’s visual generation output. For various tasks, Pisces introduces tailored visual vector lengths and employs distinct MLPs to encode conditions. This approach increases the flexibility of model design while mitigating the inference cost compared to a single encoder configuration.
@@
Current implementations of learnable query encoding can be broadly divided into two representative paradigms. The first is represented by SEED [84], which proposes a seed tokenizer that learns causal visual embeddings. Specifically, an input image is first encoded into dense token features via a BLIP-2 ViT encoder [156]. These features are then concatenated with a set of learnable query tokens and processed by a causal Q-Former to produce causal visual embeddings. This design is trained using both image-text contrastive learning and image reconstruction supervision, allowing the learned embeddings to simultaneously retain low-level visual detail and capture high-level semantic alignment with text. Building on this foundation, SEED-LLAMA [86] and SEED-X [85] enhance the model’s capacity by replacing the OPT backbone [379] with a stronger LLaMA2 model [288] and upgrading the decoder to UnCLIP-SD [244] or SDXL [224], leading to improved performance in both understanding and generation tasks. The second approach, introduced by MetaQueries [215], provides a simplified version of learnable query encoding. Here, image features are extracted via a frozen SigLIP encoder [371], which are then concatenated with learnable query tokens and directly passed through a frozen vision-language backbone such as LLaVA [33] or Qwen2.5-VL [13]. The output causal embeddings are used as conditioning inputs for a diffusion-based image decoder, enabling high-quality image generation. Because the backbone is kept frozen, the vision-language understanding capabilities remain consistent with the underlying pretrained models, offering a lightweight yet effective solution for multimodal generation. Open-Uni [323] refines the architecture of MetaQueries by utilizing solely learnable queries and a lightweight connector between a MLLM and a diffusion model, facilitating cohesive multimodal understanding and generation. Open-Uni demonstrates that the connetor between the MLLM visual understanding component and the diffusion-based visual generation component can be minimal in complexity, exemplified by a configuration comprising merely six Transformer layers. Nexus-Gen [376] and Ming-Lite-Uni [126] follow the MetaQueries paradigm, but with notable advancements to further enhance multimodal generation. Nexus-Gen [376] introduces a more powerful diffusion decoder, FLUX-1.dev, which significantly improves the generation quality. This approach allows the model to better capture the intricate details and high-fidelity features necessary for complex image generation tasks. On the other hand, Ming-Lite-Uni [126] takes a different route by introducing a highly capable MLLM model, M2-omini [100], for enhanced vision-language interaction. This model performs advanced vision-language conditioning to generate the conditioned image embeddings, ensuring a more semantically aligned representation. In addition, Ming-Lite-Uni fine-tunes its diffusion model by incorporating multi-scale learnable tokens, which facilitate improved semantic alignment across various visual scales. The multi-scale representation alignment mechanism enhances the model’s ability to generate detailed and contextually rich images from textual prompts, addressing challenges such as resolution mismatches and semantic inconsistencies. This innovative approach makes Ming-Lite-Uni a powerful tool for multimodal understanding and generation, pushing the boundaries of current methods in both flexibility and performance. UniLIP [274] incrementally incorporates reconstruction ability into CLIP through the self-distillation, then employs a learnable query along with the hidden state of the last layer of the MLLM as combined conditions. This framework is demonstrated to optimize the abundant information for visual editing. To exploit the hierarchical representations within the MLLM’s intermediate layers, TBAC-UniImage [338] applies learnable queries at multiple layers instead of the last layer. UniPic 2.0 [317] integrates SD3.5-Medium with Qwen2.5VL-7B via the MetaQuery strategy [215] and proposes a Progressive Dual-Task Reinforcement strategy for strengthening both tasks. To sum up, these learnable query-based designs share a common strength: they provide adaptive, compact, and semantically enriched representations that support both efficient image understanding and high-quality generation. By focusing on task-driven token extraction, such models offer a flexible and extensible alternative to traditional visual tokenizers, especially in unified multimodal frameworks.